OpenStudy (elvisiscool69):

solve the system by substitution . 2x+y=-11 3x-4y=11\ a.(3,5) b.(-5,-3) c.(-3,-5) d.(5,3)

OpenStudy (jtug6):

It's C. Would you like me to show you the work?

OpenStudy (elvisiscool69):

yes please

OpenStudy (elvisiscool69):

@jtug6

OpenStudy (rose4825):

Rearrange the first equation to look like this: 2x+y=-11 y=-2x-11 Then plug in the y value in the second equation: 3x-4y=11 3x - 4 (-2x - 11) = 11 3x + 8x + 44 = 11 Combine like terms: 11x + 44 = 11 Subtract 44 from each side: 11x + 44 (-44) = 11 (-44) 11x = -33 Divide 11 from each side: 11x/11 = -33/11 x = -3 Now that you have your x value (-3), you can solve for y using any of the two equations. I chose 2x + y = = -11 Plug in your x value: 2 (-3) + y = -11 -6 + y = -11 Add 6 to each side: -6 (+6) + y = -11 (+6) y = -5 Your answer would be (-3, -5)

OpenStudy (jtug6):

Oops! I'm sorry was busy rose is 100% correct. Ty rose for showing fan'd and medal'd

OpenStudy (jtug6):

lol.

OpenStudy (elvisiscool69):

rose is SO SMART!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!